OpenStudy (anonymous):

Factor and Solve x^2 – 15x + 50=0

OpenStudy (anonymous):

2. What is the Y intercept of x^2 – 15x + 50=0 3. X^2 + 6x + 12 = (x+3) True False 4. Which of the following would be the BEST problem to use the complete the square method? X ^2 + 9x + 20 x^2 - 49 x^2 -3x + 31 x^2 + 8x – 11 5. Which of the following can be factored? (Points : 2) x^2 + 10x + 20 x^2 + 10x + 21 x^2 + 10x + 30 A, B, and C can all be factored